Consider an economy that produces and consumes Bread and Automobile. Data for two different years 2005 and 2010 is given in the following table.


Year. 2005. 2010


Price of Automobiles. 5000$. 6000$


Price of a loaf of


Bread. 10$. 20$


Number of Automobiles


Produced 100$. 120$


Number of loaves


Of bread produced. 500000$. 400000$


* Using the year 2005 as a base year ,


a. Calculate the nominal and real GDP of


2010


b. Find the value of GDP Deflator for the


year 2010 and interpret your result.


c. Calculate the inflation rate in 2010.

1
Expert's answer
2021-12-30T11:33:11-0500

Solution:


a) The nominal GDP of 2010 is:


Yn = 6,000×120 + 20×400,000 = 8,720,000.


The real GDP of 2010 is:


Yn = 5,000×120 + 10×400,000 = 4,600,000.


b) The value of GDP Deflator for the year 2010 is:


Deflator = "\frac{8,720,000}{4,600,000\times100} =189.56"


So, the real GDP is almost twice lower than the nominal GDP.


c) The inflation rate in 2010 comparing to 2005 was 189.56 - 100 = 89.56%.